Larry Nelson holds 1,000 shares of General Electrics (GE) common stock. The annual stockholder meeting is being held soon, but as a minor shareholder, Larry doesnt plan to attend. Larry did not sell his shares but gave his voting rights to the management group running General Electric (GE). Larry must have signed a gives the management group control over his shares. that Larry also holds 2,000 shares of common stock in a company that only has 20,000 shares outstanding. The companys stock currently is valued at $47.00 per share. The company needs to raise new capital to invest in production. The company is looking to issue 5,000 new shares at a price of $37.60 per share. Larry worries about the value of his investment. Larrys current investment in the company is additional purchase, Larrys investment will be worth If the company issues new shares and Larry makes no This scenario is an example of Larry could be protected if the firms corporate charter includes a provision If Larry exercises the provisions in the corporate charter to protect his stake, his investment value in the firm will become
Answer #1

1) Larry must have signed a Proxy agreement(or just proxy).

In a proxy, the member of a company assigns his/her right to vote in their absence.

2) Larry currently holds 2,000 shares of 20,000 shares and the stock is currently priced at $47, hence the value of his current investment is :

=No. of shares held by Larry x Current share price

=2000*47

=$94,000

If the company issues more shares and larry does not make the additional purchase then the situation would be:

New Value of the company

= (Old shares * market price) + ( New issue * new issue price)

=(20,000*$47) + (5,000*$37.6)

=$1,128,000

Total number of shares after the additional issue is = 20,000 + 5,000 = 25,000 shares

Share price after the issue is : Total Market value / Total shares = $1,128,000 / 25,000 shares = $45.12

If Larry makes no purchase then the value of his investment will be:

= New market price * Larry's total shares

= $45.12 * 2000

=$90,240

3) The scenario is an example of dilution where the owner's number of shares are left same as before but the percentage of ownership is reduced.

Larry can be protected if the firm's charter includes a pre-emptive provision. This will give Larry the first right to purchase the additional number of shares and only if he declines then only the shares are offered to the general public.

4) If Larry exercises the provisions in the corporate charter to protect his stake, his investment value in the firm will become:

Old value = $94,000

10% of 5000 shares = 500 shares(Larry owns 10% of company)

value of new shares = 500 shares * $37.6 = $18,800

His value of investment = $94,000+$18,800 = $112,800